# Boosting Text-to-Chart Retrieval with Semantic Insights

### v2

@@ 3. Semantic Insights Synthesis Pipeline
Manually annotating charts to generate semantic insights is time-consuming and inefficient. To address this, we propose a training data development pipeline that automatically synthesizes semantic insights from raw data, ensuring the data’s effectiveness in supporting the training of text-to-chart retrieval models. This synthesized training data should enable the model to perform well across various benchmarks and improve its generalization ability.
Figure 2 shows our proposed pipeline, which consists of three key steps: data collection, automatic data visualization, and semantic insights generation based on the metadata (Figure 2: ❷–❺).
3.1. Data Collection and Visualization
@@
Step-1: Data Collection. The first step in generating high-quality semantic insights is selecting a reliable source of datasets. To meet the above criteria, we select datasets from Kaggle, which offers several advantages: many datasets are derived from real-world applications or competitions. In addition, Kaggle provides a scoring system for datasets, with those receiving a score of 10 typically offering comprehensive content and high data quality.
@@
Step-2: Automatic Data Visualization. The second step is to create meaningful charts from the collected 8,860 tables. Specifically, we used DeepEye (Luo et al., 2018a), an automatic data visualization system, to generate a variety of charts, including bar charts, pie charts, line charts, scatter plots, grouped line charts, stacked bar charts, and grouped bar charts, covering widely-used chart types (Luo et al., 2022a). In our implementation, DeepEye automatically visualized these 8,860 tables, resulting in a total of 69,166 generated charts. For each visualization, we also derived its chart specification in JSON format, which includes details like axis labels, chart titles, and data points. To enhance the visual diversity of the charts, we re-rendered them using three popular visualization libraries: Matplotlib, Plotly, and Vega-Lite. For each chart, one of these libraries was randomly selected for re-rendering (see Figure 2: ❶).
We further enriched the charts by randomizing various visualization parameters. For example, in line charts, we varied the line style between solid and dashed; in scatter plots, we alternated marker shapes (circular vs. diamond). Additionally, we introduced further variations, such as transforming a line chart into an area chart or converting a pie chart into a donut chart by adjusting a single parameter. These adjustments significantly increased the diversity of the generated charts.
@@
Note that DeepEye(Luo et al., 2018a) provides the metadata for each chart, which can be derived from its chart specification. We also store this metadata in the source code of the visualization, as shown in Figure 2. This metadata is essential for deriving semantic insights in Step 3. In total, we generated 69,166 charts using Matplotlib, Plotly, or Vega-Lite, along with their corresponding metadata.
3.2. Semantic Insights Synthesis
@@
The third step is to generate detailed chart semantic insights based on the metadata of the charts. To comprehensively interpret the charts, we design three levels of semantic insights: from the apparent visual patterns observed in the chart, to data-driven statistical insights that uncover trends, comparisons, and relationships, to task-oriented insights that provide context on how the chart can be applied in real-world scenarios for decision-making and resource allocation. This layered approach allows for a holistic interpretation of each chart, supporting diverse user queries and enhancing the retrieval process. As illustrated in Figure 2 (❷–❺), the semantic insights $I$ are computed through the following steps:
@@
Visual-oriented Insight Generation (Figure 2-❷). In this step, we generate visual-oriented insights by analyzing the chart’s visual patterns and trends. This insight summarizes the chart’s overall visual presentation, such as identifying significant trends, distributions, and patterns in the data. For instance, we would identify whether the chart shows a clear upward or downward trend, highlighting its most significant visual characteristics.
@@
Task-oriented Insight Generation (Figure 2-❸). Next, we use both the metadata and the visual insights to derive task-oriented insights. This step contextualizes the chart for practical applications, helping users understand how the chart can be used in real-world scenarios. For example, a chart depicting healthcare data might be interpreted in terms of how it could inform resource allocation or support decision-making in healthcare planning.
@@
Statistics-oriented Insight Generation (Figure 2-❺). First, we apply a set of 10 statistical analysis tasks (Wu et al., 2024) on the chart’s metadata (e.g., the $X/Y$-axes), as shown in Figure 2-❹. These tasks help identify statistical properties such as trends, correlations, and anomalies within the data. In the second step, we generate statistics-oriented insights by summarizing this statistical information. These insights provide an in-depth understanding of the chart’s data, focusing on aspects such as correlations, distributions, outliers, and other statistical relationships. This provides a deeper understanding of the data patterns and their significance.
@@
Figure 3. Evaluation of caption quality on four aspects
@@
Figure 4. (a) ChartFinder is trained using both chart images and their corresponding semantic insights. (b) ChartFinder takes a query and a set of chart images as input, retrieving the most relevant charts based on embedding similarity.
3.3. The Quality of Semantic Insights
After generating the semantic insights, we validated their quality through a crowdsourcing experiment. We randomly selected 50 captions and invited 100 crowd workers to rate them across four dimensions: completeness, consistency, diversity, and readability, using a 5-point scale (from Strongly Disagree to Strongly Agree).
The results are shown in Figure 3. Specifically, the majority of crowd workers rated the captions highly. Specifically, for completeness, 38% agreed and 37.1% strongly agreed; for consistency, 33.8% agreed and 41.3% strongly agreed; for diversity, 34.7% agreed and 40.2% strongly agreed; and for readability, 30.4% agreed and 49.8% strongly agreed. Overall, less than 5% of workers rated the semantic insights as Strongly Disagree, confirming their high quality.
These high-quality captions are used to train our text-to-chart retrieval model. If our model performs well and generalizes effectively on other text-to-chart retrieval benchmarks, it further validates the effectiveness of the training data generated by our pipeline.
4. ChartFinder: A Text-to-Chart Retrieval Model
In this section, we first describe the architecture of ChartFinder, followed by the training details. Finally, we outline how the model performs text-to-chart retrieval with ChartFinder.
4.1. Model Architecture
The design of ChartFinder is driven by the need to effectively retrieve relevant charts based on a user’s text query. This is a cross-modal retrieval task, where the goal is to align the textual queries with the charts’ visual and semantic features. Specifically, the challenge is to bridge the two modalities: the chart images (visual content) and the corresponding semantic insights (textual content). Our model should learn a shared embedding space where both visual and textual information can be effectively aligned.
To achieve this, we adopt a CLIP-based framework (Radford et al., 2021), which has demonstrated success in cross-modal tasks by aligning visual features with natural language supervision (Lin et al., 2014; Plummer et al., 2015; Chen et al., 2025). Since the semantic insights associated with charts can be long and detailed, we utilize Long-CLIP (Zhang et al., 2025) to process these insights. Unlike the original CLIP model, which is limited to encoding short captions (up to 77 tokens), Long-CLIP extends the token limit to 248 tokens, allowing it to handle the longer and more complex textual descriptions required for charts. For encoding the chart images, we use the Vision Transformer (ViT) (Dosovitskiy, 2020), which produces dense visual embeddings. ViT has shown strong performance in various image-related tasks, making it ideal for encoding the visual features of charts.
Together, the ViT-based image encoder and the Long-CLIP text encoder form the core of ChartFinder. To further explore the model’s scalability, we introduce two variants: ChartFinder and ChartFinder-B. Both versions utilize the same Long-CLIP text encoder but differ in their visual encoder. ChartFinder employs the larger ViT-L/14, while ChartFinder-B uses the smaller ViT-B/16.
4.2. Training Details
As depicted in Figure 4(a), for each chart $C_{i}$, we associate three distinct semantic insights: $I^{v}_{i}$ (visual-oriented insights), $I^{t}_{i}$ (task-oriented insights), and $I^{s}_{i}$ (statistical-oriented insights). The objective is for the model to learn to match the chart $C_{i}$ with all three semantic insights in a unified embedding space. To achieve this, we experimented with two different strategies for training:
One approach was to maintain a similarity matrix for the three different insights during training and compute the average loss across them. However, experiments showed that this method was ineffective in learning meaningful representations.
After further analysis, we decided to flatten the semantic insights corresponding to all charts and then shuffle them, as illustrated in Figure 4(a). This approach ensures that the model learns to match each chart with its corresponding semantic insights without the interference of rigid similarity constraints. Subsequently, we apply contrastive learning based on the CLIP framework to optimize the model. The contrastive loss function is defined as:
$L_{i}=-\log\frac{\exp^{(sim(C_{i},I_{i})/\tau)}}{\sum^{N}_{j=1}\exp^{(sim(C_{i},I_{j})/\tau)}},\quad I_{i},I_{j}\in[I^{v},I^{t},I^{s}]$
We trained ChartFinder using 4 A800 GPUs with a learning rate of $61e-$ 6 for 20 epochs. For the base version of the model, we set the batch size to 512, while for the larger version, we reduced the batch size to 256 due to the increased memory usage associated with the larger model.
4.3. Text-to-Chart Retrieval with ChartFinder
Text-to-chart retrieval is a cross-modal task where the goal is to retrieve the most relevant charts from a repository $\mathcal{C}=\{C_{1},C_{2},\dots,C_{n}\}$ based on a textual query $Q$ provided by the user. As shown in Figure 4(b), this process involves encoding both the text query and the chart images using text and image encoders, respectively.
Let $\vec{Q}$ represent the embedding of the query, and $\vec{\mathcal{C}}=\{\vec{C_{1}},\vec{C_{2}},\dots,\vec{C_{n}}\}$ represent the set of chart embeddings, which are generated by the ChartFinder model. These embeddings capture both the visual content of the charts and the semantic insights associated with them. Once the embeddings are generated, the retrieval process involves computing the similarity scores between the query embedding $\vec{Q}$ and all chart embeddings $\vec{\mathcal{C}}$. We calculate the similarity using a suitable similarity function, such as cosine similarity: $\text{sim}(Q,C_{i})=\frac{\vec{Q}\cdot\vec{C_{i}}}{\|\vec{Q}\|\|\vec{C_{i}}\|}$.
Finally, based on these similarity scores, we retrieve the top-$k$ most relevant charts $\{C_{r_{1}},C_{r_{2}},\dots,C_{r_{k}}\}$, where $\{C_{r_{1}},C_{r_{2}},\dots,C_{r_{k}}\}$ are the charts with the highest similarity to the query $Q$. These top-$k$ charts are presented to the user as the most relevant results based on the provided textual query.
5. CRBench Benchmark
In this section, we first discuss the motivation behind the creation of a text-to-chart retrieval benchmark (Section 5.1), followed by an introduction to the benchmark curation process (Section 5.2). Finally, we provide an overview of the key characteristics of the first text-to-chart benchmark CRBench (Section 5.3).
5.1. Motivation and Design Consideration
Text-to-chart retrieval requires retrieving relevant charts from a repository based on a user’s textual query, which involves aligning both the visual features of charts and the semantic insights embedded within the chart images. Unlike traditional image-based retrieval, which focuses solely on visual content, text-to-chart retrieval must also account for the semantic context conveyed by textual descriptions.
While datasets like VisText (Tang et al., 2023) and Chart-to-Text (Kantharaj et al., 2022) exist, they primarily focus on chart captioning and description generation, rather than the full text-to-chart retrieval task. These datasets aim to generate textual descriptions of charts but do not directly address the challenge of retrieving the appropriate chart based on textual queries, which is central to our work.
As a result, there is currently no benchmark for evaluating the performance of the text-to-chart retrieval task, which involves the alignment of both textual queries and chart visual content. This gap highlights the need for a benchmark that can evaluate the performance of models on this task.
Design Consideration. To fill this gap, we aim to propose a real-world dataset designed for text-to-chart retrieval, enabling the fair comparison of retrieval models across a variety of use cases. Several key factors were considered, as outlined below:
(1) Precise and Fuzzy Queries: As discussed in Section 1, user queries in real-world BI applications can be both precise (focused on specific details) and fuzzy (more general). Our benchmark includes both types of queries to reflect real-world user behavior and better evaluate model performance across different query complexities.
(2) Diversity of Chart Types: To reflect the variety of data visualizations encountered in real-world applications, the benchmark includes a broad range of chart types, testing model performance across various chart structures.
(3) Real-world Charts: The charts included in the benchmark should be sourced from real-world applications, ensuring their applicability in addressing actual user needs and scenarios.
The prompt for GPT-4o to generate queries. System Prompt: You are a user searching for visualizations in a database. You will see 5 visualizations, but you should ONLY focus on the FIRST image when generating queries. The other four images are very similar to the first one, make sure your queries CAN NOT match any of the other four images. Generate two meaningful queries (10-15 words each): 1. Precise Query: Generate a query about specific content in the FIRST visualization that includes: - Meaningful combination of axis labels/categories - Important data values or time ranges - Key categories or measurements - If image contains time, you should include the time range in the query because it is important for the user to search for the specific time range and distinguish from other similar visualizations. - When cover time range, do not direct use the initial time range in the chart. For example, if the initial time range in the x-axis is 2012-2019, you can use 2014 - 2018 or 2013 - 2016. Make sure the time range in your query is included in the initial time range. - You don’t need to cover very specific time range like specific months or days unless you need this to distinguish from other similar visualizations. - Do not use comma in your query. - Do not directly use the initial text in the image as the query, always find similar words or phrases to describe the content of the image. For example, if the title contains ”weekly”, you should use every 7 days in the query. Example: ”global temperature change 1990-2020 in the United States” (good, time and location are included) Bad example: ”global temperature change” (bad, no time and location to distinguish from other similar visualizations). 2. Fuzzy Query: Generate a query about the visualization purpose that includes: - Line charts: trend analysis or comparison over time - Bar charts: value comparison or ranking purpose - Pie/stacked charts: distribution analysis or proportion comparison - Scatter plots: correlation or pattern analysis Example: ”annual economic growth comparison between countries” (enough information to distinguish from other similar visualizations) Bad example: ”growth comparison” (too little information) Format your response as: { ”Precise query”: ”¡10-15 word query about content¿”, ”Fuzzy query”: ”¡10-15 word query about purpose¿” } Remember: - ONLY consider the FIRST image - Include relevant context - Be specific and meaningful - Articles (a, an, the) and prepositions count as words
Figure 5. The prompt for GPT-4o to generate precise queries and fuzzy queries.
5.2. Benchmark Construction
To build a benchmark that meets the aforementioned considerations, we first collected a large set of charts from real-world BI applications. Next, we used a chart embedding model to map the relevant charts to a given text query. We then employed crowdsourcing to perform preliminary annotations based on these text-chart pairs. Next, GPT-4o was used to generate text queries based on the predefined templates. Finally, we conducted a second round of crowdsourcing to validate the accuracy and relevance of the model-generated text queries.
5.2.1. Charts Collection from Real-world Applications.
To ensure CRBench reflects real-world chart retrieval scenarios, we collected charts from reliable sources widely used in business intelligence and data analysis: (i) Tableau: A leading data visualization platform for charting and business analytics (Tableau, 2003). (ii) Statista: A platform providing statistical data across various sectors (statista, 2007). (iii) Pew Research Center: A non-profit organization focusing on social science research (Research, 2004). (iv) Our World in Data: An open-access platform researching global challenges through charts on topics like economics, health, and society (OWID, 2011).
From these sources, we gathered 21,862 charts covering various domains, including finance, business, e-commerce, technology, and more. These fields are closely related to BI applications.
5.2.2. Text Query Generation and Annotation
After collecting charts that meet the BI scenario, our next step is to generate and annotate text queries based on the chart repository.
Step 1: Similar Charts Selection. In the real world, for example, when searching for charts in a BI system, the number of charts should be much larger than the number of queries. Moreover, each query will correspond to many similar charts. Based on this setting, the first step in annotating a query is to find similar charts. Calculating the similarity of rasterized charts is rather difficult. Thus, we encode the charts into different embeddings and compare their representation similarity.
We use a pre-trained ViT model to encode chart $C_{i}$: $\vec{C}_{i}=ViT(C_{i}),i\in[1,21862]$. $\vec{C}_{i}$ is the embedding vector of the $i$-th chart, with dimension 1 $\times$ 768. Next, we concatenate all the single embeddings and have all the tensors $T$: $T=Concat[\vec{C}_{1};\vec{C}_{2};...;\vec{C}_{21862}]$. $T$ shapes like 21862 $\times$ 768. After we get all the embedding tensors, we calculate the similarity matrices between these tensors with cosine similarity: $sim=\cos(T,T^{T})$. Based on the similarity matrix, we can now easily get the similarity score of candidate $C_{j}$ and the target $C_{i}$: $S_{ij}=sim(C_{i},C_{j}),i,j\in[1,21862]$. Afterward, we will set $\theta$ as the similarity threshold. Only if $S_{ik}>\theta,\forall k\in[j_{1},j_{2},j_{3},j_{4}]$. Then, we combine all four candidate charts and the target chart as one chart group: $Group_{a}=(C_{a1},C_{a2},C_{a3},C_{a4},C_{a5}),a\in[1,274]$.
In this step, we set the threshold as 0.90 and produced 247 groups of similar charts.
Step 2: Text Query Annotation by Crowdsourcing. We collect text queries through the Appen (Ltd, 2009) crowdsourcing platform based on groups of similar charts collected in Step 1. In this step, we ask workers to write both precise and fuzzy queries for each chart. To ensure the relevance of the text queries, we provide each worker not only with the chart in question but also with the four most similar charts to the first chart. These additional charts serve as references, ensuring that the queries are contextually aligned with the chart’s content and not influenced by unrelated charts.
In the first round of crowdsourcing experiments, we assign five workers to annotate queries for each grouped chart. However, due to the inherent variability of crowdsourcing and the complexity of the task, we found that some queries were completely unrelated to the chart. Specifically, we observed that 2 or 3 responses in each group of submissions were irrelevant.
Step 3: Text Query Generation by GPT-4o. Building on the query patterns identified in the first 50 crowdsourcing experiments, we used them to generate high-quality prompt words for GPT-4o. The initial crowdsourcing data was found to have significant variability and low quality, which made it less reliable for further use. To improve the query quality, we turned to GPT-4o to generate queries.
Following the method outlined in Step 2, we show GPT-4o five charts at a time and instruct it to generate two types of queries. Specifically, we use a carefully designed prompt, which is shown in Figure 5, to guide GPT-4o to generate precise queries and fuzzy queries for only one of the five similar charts. We specifically ensure that the precise query contains enough details and constraints to make it accurately match only the target chart, not other charts. This method ensures that the queries generated are tailored to each chart’s characteristics and use case, addressing the challenges of crowdsourcing quality and providing more relevant and precise queries for chart retrieval.
Table 1. Comparison of Existing Benchmarks. The statistics for Chart-to-Text and VisText are from the respective test sets. The column “C/Q” indicates the ratio for charts (C) and text queries (Q). The column “Dist. of Query Len.” Shows the distribution of lengths of text queries in characters. The column “CT” means the types of charts.
Datasets CT #-Charts #-Queries C/Q Difficulty Dist. of Query Len. Samples of Similar Charts
VisText (Tang et al., 2023) 3 882 1764 0.5 Easy [Uncaptioned image] [Uncaptioned image]
Chart-to-Text (Kantharaj et al., 2022) 6 1393 1393 1 Easy [Uncaptioned image] [Uncaptioned image]
CRBench (ours) 11 21,862 326 67 Hard [Uncaptioned image] [Uncaptioned image]
Step 4: Text Query Annotation by Crowdsourcing. After generating a complete set of 247 precise and fuzzy queries, we noticed that, despite clear instructions to ensure the generated queries were relevant to the first chart and distinct from the other charts in the set, some queries still applied to multiple charts. To address this issue, we conducted a second round of crowdsourcing experiments.
In this round, we showed the generated queries along with their corresponding five charts to nine workers. We asked the workers to select the chart that they believed best matched the query. If five or more workers selected the same chart as the target, we considered the query to be unambiguous. After this crowdsourcing screening process, we retained 195 precise queries and 131 fuzzy queries.
5.3. CRBench Overview
In this section, we compare CRBench with two existing benchmarks, Chart-to-Text (Kantharaj et al., 2022) and VisText (Tang et al., 2023), both of which focus on chart-related tasks like caption generation and description. While these benchmarks are useful for evaluating chart captioning and description generation, they do not address the text-to-chart retrieval task. In our study, we repurpose the captions in these benchmarks as queries for retrieval tasks, enabling us to compare their effectiveness with CRBench in retrieving relevant charts based on text queries. Specifically, in VisText, each chart is paired with two types of captions: L1 captions (e.g., basic properties) and L2+L3 captions (e.g., trends and statistics). This results in a chart-to-query ratio of 1:2. In contrast, Chart-to-Text pairs each chart with a single caption, resulting in a 1:1 ratio. In CRBench, however, the ratio is much higher (67:1), meaning the number of charts significantly exceeds the number of queries, which is more representative of real-world scenarios where charts vastly outnumber queries.
Moreover, we compared examples of similar charts across the benchmarks. In VisText, similar charts show strong visual resemblance but often differ significantly in semantic content. In Chart-to-Text, charts deemed similar tend to align well semantically but differ visually. On the other hand, CRBench ensures that similar charts are closely aligned both visually and semantically, resulting in a more holistic and accurate comparison.

### v4

@@ 4 Semantic Insights Synthesis Pipeline
As we mentioned in the Introduction, we need high-quality caption data for model training to retrieve correct charts. However, manually annotating charts to generate semantic insights is time-consuming and inefficient. To address this, we propose a training data development pipeline that automatically synthesizes semantic insights from raw data, ensuring the data’s effectiveness in supporting the training of text-to-chart retrieval models. Figure 2 shows our proposed pipeline, which consists of three key steps: data collection, automatic data visualization, and semantic insights generation based on the metadata (Figure 2: ❷–❺).
4.1 Data Collection and Visualization
@@
Step-1: Data Collection. The first step in generating high-quality semantic insights is selecting a reliable source of datasets. To meet the above criteria, we select datasets from Kaggle, which offers several advantages: first, almost all the datasets are derived from real-world applications or competitions. Second, Kaggle provides a scoring system for datasets, with those receiving a score of 10 typically offering comprehensive content and high data quality.
@@
Step-2: Automatic Data Visualization. Next, we used DeepEye Luo et al. (2018a), an automatic data visualization system, to generate a variety of charts, including bar charts, pie charts, line charts, scatter plots, grouped line charts, stacked bar charts, and grouped bar charts, covering widely-used chart types Luo et al. (2022a). In our implementation, DeepEye generated 69,166 charts based on the 8,860 tables. For each visualization, we also derived its chart metadata in JSON format, which includes details like axis labels, chart titles, and data points. To enhance the visual diversity of the charts, we re-rendered them by randomly using a visualization libraries from Matplotlib, Plotly, and Vega-Lite (see Figure 2: ❶).
@@
We further enhance the visual diversity of generated charts by systematically randomizing various visual parameters. For example, adjusting line type, marker shape, or switching between sub-types such as pie and donut charts. We also store the metadata produced by DeepEye Luo et al. (2018a) in the source code of the visualization, as shown in Figure 2. This metadata is essential for deriving semantic insights in Step 3. In total, we generated 69,166 charts using Matplotlib, Plotly, or Vega-Lite, along with their corresponding metadata.
4.2 Semantic Insights Synthesis
@@
The third step is to generate detailed chart semantic insights based on the metadata of the charts. To comprehensively interpret the charts, we design three levels of semantic insights: (1) The apparent visual patterns observed in the chart. (2) Data-driven statistical insights that uncover trends, comparisons, and relationships. (3) Task-oriented insights that provide context on how the chart can be applied in real-world scenarios for decision-making and resource allocation. This layered approach allows for a holistic interpretation of each chart, supporting diverse user queries and enhancing the retrieval process. As illustrated in Figure 2 (❷–❺), the semantic insights $I$ are generated through the following steps:
@@
Visual-oriented Insight Generation (Figure 2-❷). In this step, we generate visual-oriented insights by analyzing the chart’s visual patterns and trends. This insight summarizes the chart’s overall visual presentation, such as identifying significant trends, distributions, and patterns in the data.
@@
Task-oriented Insight Generation (Figure 2-❸). Next, we use both the metadata and the visual insights to derive task-oriented insights. This step contextualizes the chart for practical applications, helping users understand how the chart can be used in real-world scenarios.
@@
Statistics-oriented Insight Generation (Figure 2-❺). First, we apply a set of 10 statistical analysis tasks Wu et al. (2024) on the chart’s metadata (e.g., the $X/Y$-axes), as shown in Figure 2-❹. These tasks help identify statistical properties such as trends, correlations, and anomalies within the data. In the second step of our pipeline, we generate statistics-oriented insights by summarizing this statistical information.
This provides a deeper understanding of the data patterns and their significance.
@@
Figure 3: Evaluation of caption quality on four aspects
@@
Figure 4: (a) ChartFinder is trained using both chart images and their corresponding semantic insights. (b) ChartFinder takes a query and a set of chart images as input, retrieving the most relevant charts based on embedding similarity.
4.3 The Quality of Semantic Insights
To verify the quality of the semantic insights generated, we conducted a crowdsourcing experiment. We randomly selected 50 semantic insights and invited 100 crowdworkers to rate them on four dimensions: Completeness, Consistency, Diversity, and Readability, using a 5-point scale (from “strongly disagree" to “strongly agree"). As shown in Figure 3, the evaluation results are very positive. The total number of positive ratings of “agree" and “strongly agree" exceeded 74% in all four dimensions, especially readability, which received more than 80% approval. This result fully confirms the high quality of our synthesized insights. These high-quality data are the cornerstone of our subsequent training of text-graph retrieval models. The good generalization ability of the model in other benchmarks also further verifies the effectiveness of our data generation pipeline.
5 Experiments
In this section, we conduct a set of experiments to evaluate whether the synthesized semantic insights can enhance text-to-chart retrieval performance across multiple benchmarks and settings. We also assess whether the ChartFinder model achieves state-of-the-art performance against all baselines. The following experiments are performed:
Exp-1: How does the overall performance of ChartFinder compare with other baselines on CRBench?
Exp-2: How does ChartFinder perform when the text query becomes a long caption?
Exp-3: Ablation study on the combination of synthesized semantic insights.
Exp-4: The generalizability of synthesized semantic insights.
Exp-5: The impact of center crop strategy on text-to-chart retrieval. In the main body of the paper, we will focus on the detailed analysis of Exp-1. Due to page limitations, the comprehensive results and discussions for Exp-2 through Exp-5 are provided in the Appendix A.